The length of a rectangular frame is 12 inches more than twice the width. The area inside the frame is 1040 square inches. Find the dimension of the frame.

Answer by JimboP1977(311)

Length = L
Width = W
L+12 = 2W
L x W = 1040
(2W-12) x W = 1040
2W^2 -12W = 1040
W^2 - 6W = 520
(W-3)^2 - 9 = 520
W = 529^(1/2) + 3
W = 26
L = 40